The absolute visual magnitude of a star is 6.5 and for its temperature, the bolometric correction is - 0.3. Calculate the absolute bolometric magnitude and the luminosity of the star.
1
Expert's answer
2018-10-23T11:38:09-0400

First we need to obtain absolute bolometric magnitude: M_bol=M_Vis+∆_bol=6.5-0.3=6.2


Now we can compare this value to the Sun’s: M_sun-M_bol=-2.5 log_10⁡〖L_sun/L_star 〗

This gives us: L_star=L_sun 〖10〗^0.4(M_sun-M_bol ) =3*〖10〗^28*〖10〗^0.4(4.74-6.2) W/m^2=7.8*〖10〗^27 W/m^2


So, L_star=7.8*〖10〗^27 W/m^2